OpenStudy (anonymous):

Solve the variable in CAPS a= B -2c

OpenStudy (anonymous):

Solve for B? So, add 2c on both sides. We get a + 2c = B - 2c + 2c B = a + 2c.

OpenStudy (anonymous):

Okay , Thanks (: !!

OpenStudy (anonymous):

No problem :-)